# Elimination distance to bounded degree on planar graphs

### v3

@@ 1 Introduction
For $k,d\in\mathbb{N}$, we denote by $\mathscr{C}_{k,d}$ the class of all graphs that have elimination distance at most $k$ to $\mathscr{C}_{d}$. It is easy to see that for every fixed $k$ and $d$ we can formulate the property that a graph is in $\mathscr{C}_{k,d}$ by a sentence in monadic second-order logic (MSO). By the famous theorem of Courcelle [9] we can test every MSO-property $\varphi$ in time $f(|\varphi|,t)\cdot n$ on every $n$-vertex graph of treewidth $t$ for some computable function $f$. Hence, we can decide for every $n$-vertex graph $G$ of treewidth $t$ whether $G\in\mathscr{C}_{k,d}$ in time $f(k,d,t)\cdot n$ for some computable function $f$. However, for $d\geqslant 3$ already the class $\mathscr{C}_{d}$ has unbounded treewidth, and so the same holds for $\mathscr{C}_{k,d}$ for all values of $k$. Thus, Courcelle’s Theorem cannot be directly applied to derive fixed-parameter tractability of the problem in full generality.
@@
Theorem 1 (Main result).
@@
We first presented Theorem 1 in a conference paper [20]. In this journal version, the presentation of the proof has been vastly improved and simplified. While the main result is the same, some observations enabled us to remove unnecessary complicated notion and to simplify the proof.
@@ 2 Preliminaries
A partial order on a set $V$ is a binary relation $\leqslant$ on $V$ that is reflexive, anti-symmetric and transitive. A set $W\subseteq V$ is a chain if it is totally ordered by $\leqslant$. If $\leqslant$ is a partial order on $V$, and for every element $v\in V$ the set $V_{\leqslant v}\coloneqq\{u\in V\mid u\leqslant v\}$ is a chain, then $\leqslant$ is a tree order. Note that the covering relation of a tree order is not necessarily a tree, but may be a forest. An elimination order on a graph $G$ is a tree order $\leqslant$ on $V(G)$ such that for every edge $\{u,v\}\in E(G)$ we have either $u\leqslant v$ or $v\leqslant u$. The depth of a vertex $v$ in a tree order $\leqslant$ is the size the set $V_{<v}\coloneqq\{u\in V\mid u<v\}$. The depth of a tree order $\leqslant$ is maximal depth among all vertices.
@@
We denote by $\mathscr{C}_{d}$ the class of all graphs of maximum degree at most $d$ and by $\mathscr{C}_{k,d}$ the class of all graphs with elimination distance at most $k$ to $\mathscr{C}_{d}$. Note for instance that $\mathrm{td}(G)=k$ if and only if $G\in\mathscr{C}_{k,0}$. We write $\mathrm{ed}_{d}(G)$ for $\mathrm{ed}_{\mathscr{C}_{d}}(G)$.
Definition 2 (Definition 4.2 of [6]).
A tree order $\leqslant$ on G is an elimination order to degree $d$ if for every $v\in V(G)$ the set $S_{v}\coloneqq\{u\in G\mid\{u,v\}\in E(G),u\not\leqslant v$ and $v\not\leqslant u\}$ satisfies either:
@@
    $v$ is $\leqslant$-maximal, $|S_{v}|\leqslant d$, and for all $u\in S_{v}$, we have $\{w\mid w<u\}=\{w\mid w<v\}$.
@@
Proposition 3 (Proposition 4.3 in [6]).
A graph $G$ satisfies $\mathrm{ed}_{d}(G)\leqslant k$ if, and only if, there exists an elimination order to degree $d$ of depth $k$ for $G$.
@@
Lemma 4.
For every graph $G$ and elimination order $\leqslant$ to degree $d$ for $G$, we can compute in quadratic time an elimination order $\preceq$ to degree $d$ for $G$, with depth not larger than the depth of $\leqslant$, and with the additional property that for every $v\in V(G)$, if $C,C^{\prime}$ are distinct connected components of $G-V_{\preceq v}$ (or of $G$), then the vertices of $C$ and $C^{\prime}$ are incomparable with respect to $\preceq$.
Let $G$ be a graph. A graph $H$ with vertex set $\{v_{1},\ldots,v_{n}\}$ is a minor of $G$, written $H\preceq G$, if there are connected and pairwise vertex disjoint subgraphs $H_{1},\ldots,H_{n}\subseteq G$ such that if $\{v_{i},v_{j}\}\in E(H)$, then there are $w_{i}\in V(H_{i})$ and $w_{j}\in V(H_{j})$ such that $\{w_{i},w_{j}\}\in E(G)$. We call the subgraph $H_{i}$ the branch set of the vertex $v_{i}$ in $G$. If all $H_{i}$ have radius at most $r$, then we say that $H$ is a depth-$r$ minor of $G$, in symbols $H\preceq_{r}G$. If $G$ is a graph and =$\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ is a set of pairwise vertex disjoint subgraphs of $G$, then the graph $H$ with vertex set $\{v_{1},\ldots,v_{n}\}$ and edges $\{v_{i},v_{j}\}\in E(H)$ if and only if there is an edge between a vertex of $H_{i}$ and a vertex of $H_{j}$ in $G$, the minor induced by $\mathcal{H}$. If $\bigcup_{1\leqslant i\leqslant n}V(H_{i})=V(G)$, we call $\mathcal{H}$ a minor model of $H$ that subsumes all vertices of $G$.
We denote by $K_{t}$ the complete graph on $t$ vertices. We denote by $G_{m\times n}$ the grid with $m$ rows and $n$ columns, that is, the graph with vertex set $\{v_{i,j}\mid 1\leqslant i\leqslant m,1\leqslant j\leqslant n\}$ and edges $\{v_{i,j},v_{i^{\prime},j^{\prime}}\}$ for $|i-i^{\prime}|+|j-j^{\prime}|=1$.
@@
Theorem 5 (Excluded Grid Theorem).
There exists a function $g$ such that for every integer $t\geqslant 1$, every graph of treewidth at least $g(t)$ contains the $t\times t$ grid as a minor. Furthermore, such a grid minor can be computed in polynomial time.
@@
Theorem 6 (Planar Excluded Grid Theorem, see Theorem 7.23 of [10]).
@@
Theorem 7 (Courcelle’s Theorem [9]).
@@ 3 The proof
For $q,k\in\mathbb{N}$, a graph $G$ is $(q,k)$-unbreakable if for every separation $(A,B)$ of $G$ of order at most $k$, we have $|A|\leqslant q$ or $|B|\leqslant q$.
Lemma 8.
@@
Proof.
Let $G=G_{m\times m}$ and let $A,B$ be a separation of order at most $k$ in $G$. We show that $|A|\leqslant k^{2}$ or $|B|\leqslant k^{2}$. The statement is trivially true if $m\leqslant k$ as then $|V(G)|\leqslant k^{2}$. Hence, assume that $m>k$, and let $\overline{z}\coloneqq z_{1},\ldots,z_{k}=A\cap B$.
We call a row or a column of $G$ untouched if it does not intersect $\overline{z}$ and touched if it does intersect $\overline{z}$. As there are at least $k+1$ rows (resp. columns) there is at least one untouched row (resp. column). An untouched row (resp. column) must be contained in one part of the separation. Assume w.l.o.g. that $A$ contains an untouched row. Then $A$ contains all untouched columns and, by symmetry, all untouched rows. Therefore every element in $B$ must be part of a touched row and a touched column, and, as there are at most $k$ touched rows resp. columns, there are at most $k^{2}$ such elements. ∎
@@
Corollary 9 (of Lemma 8).
For all integers $k,m$ with $k<m$, for all connected graphs $G$ for which there exists a minor model $\mathcal{G}_{m,m}=\{H_{i,j}~{}|~{}1\leqslant i,j\leqslant m\}$ that subsumes all vertices of $G$, $k$ rounds of recursive elimination on $G$ can delete vertices in at most $k^{2}$ many different branch sets.
Definition 10.
Let $G$ be a $K_{5}$-minor-free graph and let $k,d\in\mathbb{N}$. Assume there exists a minor model $\mathcal{G}_{m,m}=\{H_{i,j}\mid 1\leqslant i,j\leqslant m\}$ (for $m\geqslant 4k+7$) that subsumes all vertices of $G$ and that induces a supergraph of the grid $G_{m,m}$. We call the branch set $H_{i,j}$ $(k,d)$-safe if $32k+3\leqslant i,j\leqslant m-2k-$ 3 and if $H_{i^{\prime},j^{\prime}}$ contains no vertex of degree at least $d+1$ (in $G$) for $|i-i^{\prime}|,|j-j^{\prime}|\leqslant 2k+2$.
Lemma 11.
Let $G$ be a $K_{5}$-minor-free graph and let $k,d\in\mathbb{N}$. Assume there exists a minor model $\mathcal{G}_{m,m}=\{H_{i,j}\mid 1\leqslant i,j\leqslant m\}$ (for $m\geqslant 4k+7$) that subsumes all vertices of $G$ and that induces a supergraph of the grid $G_{m,m}$. Assume $H_{i,j}$ is $(k,d)$-safe. Let $a\in V(H_{i,j})$, let $B\subseteq V(G)\setminus\{a\}$ with $|B|\leqslant k$ and let $x,y\in V(G)\setminus(B\cup\{a\})$ be of degree at least $d+1$. Then $x$ and $y$ are connected in $G-B$ if and only if $x$ and $y$ are connected in $G-B-a$.
Proof.
@@
We define sets $X_{\ell}$ for $10\leqslant\ell\leqslant k+$ 1 of branch sets as follows. Let $X_{0}$ be the set consisting only of $H_{i,j}$ and for $\ell\geqslant 1$ let $X_{\ell}$ be the set of all $H_{i^{\prime},j^{\prime}}$ with $2\ell-1\leqslant|i-i^{\prime}|,|j-j^{\prime}|\leqslant 2\ell$ that do not already belong to $X_{\ell-1}$. The sets $X_{\ell}$ are the borders (of thickness $2$) of the $(4\ell+1)\times(4\ell+1)$-subgrid around $H_{i,j}$. Observe that the vertices $x$ and $y$ do not belong to any of the $X_{\ell}$, as $H_{i,j}$ is $(k,d)$-safe by assumption. For $10\leqslant\ell\leqslant k+$ 1 let $Y_{\ell}$ be the subgraph of $G$ induced by the vertices of $X_{\ell}$.
We claim that for $11\leqslant\ell\leqslant k+$ 1, the sets $Y_{\ell}$ are connected sets that separate $a$ from $x$ and analogously $a$ from $y$. Clearly, the $Y_{\ell}$ are connected. Now observe that there is no edge between a vertex of $\bigcup_{0\leqslant i\leqslant\ell-1}Y_{i}$ and a vertex of $G-\bigcup_{0\leqslant i\leqslant\ell}Y_{i}$. The existence of such a connection would create a $K_{5}$ minor, see [10, Figure 7.10] or Figure 1. Hence, any path between $a$ and $x$ (or $y$) must pass through $Y_{\ell}$.
As $|B|\leqslant k$, there is one $Y_{\ell}$ with $11\leqslant\ell\leqslant k+$ 1 that does not intersect $B$. Let $u$ be the first time that $P$ visits $Y_{\ell}$ on its way from $x$ to $a$ and let $v$ be the last time that $P$ visits $Y_{\ell}$ on its way from $a$ to $y$. As $Y_{\ell}$ is connected, we can reroute the subpath between $u$ and $v$ through $Y_{\ell}$ and thereby construct a path between $x$ and $y$ in $G-B-a$. ∎
Figure 1: Construction of a $K_{5}$ minor as soon as a branch set (here $H_{1}$) connected to a branch set (here $H_{2}$) that is at distance more than 2 in the grid [10, Figure 7.10]. The blue part marks the “outer part” of the border of thickness $2$, the “inner part” decomposes into a green, yellow and gray part.
@@
Corollary 12.
Let $G$ be a $K_{5}$-minor-free graph and let $k,d\in\mathbb{N}$. Assume there exists a minor model $\mathcal{G}_{m,m}=\{H_{i,j}\mid 1\leqslant i,j\leqslant m\}$ (for $m\geqslant 4k+7$) that subsumes all vertices of $G$ and that induces a supergraph of the grid $G_{m,m}$. Assume $H_{i,j}$ is $(k,d)$-safe. Then every vertex $a\in H_{i,j}$ is irrelevant, i.e., $G-a\in\mathscr{C}_{k,d}$ if and only if $G\in\mathscr{C}_{k,d}$.
Proof.
Let $H\coloneqq G-a$. We have to prove that $H\in\mathscr{C}_{k,d}$ implies $G\in\mathscr{C}_{k,d}$. Hence, assume $H\in\mathscr{C}_{k,d}$. Let $\leqslant_{H}$ be an elimination order to degree $d$ of depth $k$ for $H$. We also assume that $\leqslant_{H}$ satisfies the property of Lemma 4, that is, for every $v\in V(G)$, if $C,C^{\prime}$ are distinct connected components of $G-V_{\leqslant_{H}v}$, then the vertices of $C$ and $C^{\prime}$ are incomparable with respect to $\leqslant_{H}$. Let $A_{0}$ be the connected component of $G$ containing $a$. Note that $A_{0}$ may break into multiple connected components in $H=G-a$. We show below that at most one of these connected components contains a vertex of degree excedeeing $d$.
For $1\leqslant i\leqslant k$, we define inductively:
  • $m_{i}$ as the unique $\leqslant_{H}$-minimal element of $A_{i-1}\setminus\{a\}$ (if it exists) such that there exists a vertex $v$ with $m_{i}\leqslant_{H}v$ and of degree at least $d+1$ in $H[A_{i-1}\setminus\{a\}]$. If there is no such element $m_{i}$, the process stops.
    Let us prove that there is at most one candidate for $m_{i}$. Assume that there are incomparable $m$ and $m^{\prime}$ satisfying these conditions. This means that there are vertices $v$ and $v^{\prime}$ of degree at least $d+1$ in $A_{i-1}\setminus\{a\}$ (hence of degree at least $d+1$ in $G$) with $m\leqslant_{H}v$ and $m^{\prime}\leqslant_{H}v^{\prime}$. Note that we have $m\not\leqslant_{H}v^{\prime}$ because $\leqslant_{H}$ is a tree order. By Lemma 4, we have that $v,v^{\prime}$ are both in $A_{i-1}\setminus\{a\}$, i.e. in the connected component of $a$ in $G-\{m_{1},\ldots,m_{i-1}\}$ (it follows also by induction that $\{m_{1},\ldots,m_{i-1}\}=V_{\leqslant_{H}m_{i-1}}$, hence we may apply the lemma). Hence, $v$ and $v^{\prime}$ are connected in $G-\{m_{1},\ldots,m_{i-1}\}$. With Lemma 11, we also have that $v$ and $v^{\prime}$ are connected in $G-\{a,m_{1},\ldots,m_{i-1}\}$.
    We take a witness path from $v$ to $v^{\prime}$. Since $m\leqslant_{H}v$ and $m\not\leqslant_{H}v^{\prime}$, this path must contain two adjacent vertices $w$, $w^{\prime}$ with $m\leqslant_{H}w$ and $m\not\leqslant_{H}w^{\prime}$. This contradicts the fact that $\leqslant_{H}$ is an elimination order satisfying the property of Lemma 4. Therefore, there is at most one possible such $m_{i}$.
  • $T_{i}:=\{v\in A_{i-1}~{}:~{}m_{i}\not\leqslant_{H}v\}$. These are the vertices that are disconnected from $A_{i}$ in $H$, but not in $G$. As explained above, they all have degree at most $d$.
@@
The processes stops after at most $k$ rounds. When the process stops, we have then defined $m_{i}$, $T_{i}$ and $A_{i}$ up to $i=\omega$, with $\omega\leqslant k$ and every element in $A_{\omega}$ has degree at most $d$ in $H[A_{\omega}]$.
We then define the new order $\leqslant_{G}$ as follows:
  • for all $x,y$ other than $a$ and that are not in any of the $T_{i}$ nor in $A_{\omega}$, we have $x\leqslant_{G}y$ if and only if $x\leqslant_{H}y$,
  • for all $x$ in $A_{\omega}\cup\bigcup\limits_{i\leqslant\omega}T_{i}$, we have $m_{i}\leqslant_{G}x$ for all $i\leqslant\omega$. Note that $a$ is in $A_{\omega}$.
Note that all the elements in $A_{\omega}$, and the $T_{i}$’s, together with $a$ are $\leqslant_{G}$-maximal.
We now prove that this new order is indeed an elimination order to degree $d$ of depth $k$ for $G$. We have that $\leqslant_{G}$ is a tree order and that it has depth at most $k$. Let us now take a vertex $b$ and study $S_{b}$. Recall the definition of $S_{b}$ from Definition 2. As we have two orders, we distinct $S^{G}_{b}$ from $S^{H}_{b}$.

### v4

@@ 1 Introduction
For $k,d\in\mathbb{N}$, we denote by $\mathscr{C}_{k,d}$ the class of all graphs that have elimination distance at most $k$ to $\mathscr{C}_{d}$. It is easy to see that for every fixed $k$ and $d$ we can formulate the property that a graph is in $\mathscr{C}_{k,d}$ by a sentence in monadic second-order logic (MSO). By the famous theorem of Courcelle [11] we can test every MSO-property $\varphi$ in time $f(|\varphi|,t)\cdot n$ on every $n$-vertex graph of treewidth $t$ for some computable function $f$. Hence, we can decide for every $n$-vertex graph $G$ of treewidth $t$ whether $G\in~{}\mathscr{C}_{k,d}$ in time $f(k,d,t)\cdot n$ for some computable function $f$. However, for $d\geq 3$ already the class $\mathscr{C}_{d}$ has unbounded treewidth, and so the same holds for $\mathscr{C}_{k,d}$ for all values of $k$. Thus, Courcelle’s Theorem cannot be directly applied to derive fixed-parameter tractability of the problem in full generality.
@@
Theorem 1.1 (Main result)
@@
We first presented Theorem 1.1 in a conference paper [17]. In this journal version, the presentation of the proof has been vastly improved and simplified. While the main result is the same, some observations enabled us to remove unnecessary complicated notion and to simplify the proof.
@@ 2 Preliminaries
A partial order on a set $V$ is a binary relation $\leq$ on $V$ that is reflexive, anti-symmetric and transitive. A set $W\subseteq V$ is a chain if it is totally ordered by $\leq$. If $\leq$ is a partial order on $V$, and for every element $v\in V$ the set $V_{\leq v}\coloneqq\{u\in V\mid u\leq v\}$ is a chain, then $\leq$ is a tree order. Note that the covering relation of a tree order is not necessarily a tree, but may be a forest. An elimination order on a graph $G$ is a tree order $\leq$ on $V(G)$ such that for every edge $\{u,v\}\in E(G)$ we have either $u\leq v$ or $v\leq u$. The depth of a vertex $v$ in a tree order $\leq$ is the size the set $V_{<v}\coloneqq\{u\in V\mid u<v\}$. The depth of a tree order $\leq$ is maximal depth among all vertices.
@@
We denote by $\mathscr{C}_{d}$ the class of all graphs of maximum degree at most $d$ and by $\mathscr{C}_{k,d}$ the class of all graphs with elimination distance at most $k$ to $\mathscr{C}_{d}$. Note for instance that $\mathrm{td}(G)=k$ if and only if $G\in~{}\mathscr{C}_{k,0}$. We write $\mathrm{ed}_{d}(G)$ for $\mathrm{ed}_{\mathscr{C}_{d}}(G)$.
Definition 2.1 (Definition 4.2 of [3])
A tree order $\leq$ on G is an elimination order to degree $d$ if for every $v\in V(G)$ the set $S_{v}\coloneqq\{u\in G\mid\{u,v\}\in E(G),u\not\leq v$ and $v\not\leq u\}$ satisfies either:
@@
    $v$ is $\leq$-maximal, $|S_{v}|\leq d$, and for all $u\in S_{v}$, we have $\{w\mid w<u\}=\{w\mid w<v\}$.
@@
Proposition 2.2 (Proposition 4.3 in [3])
A graph $G$ satisfies $\mathrm{ed}_{d}(G)\leq k$ if, and only if, there exists an elimination order to degree $d$ of depth $k$ for $G$.
@@
Lemma 2.3
For every graph $G$ and elimination order $\leq$ to degree $d$ for $G$, we can compute in quadratic time an elimination order $\preceq$ to degree $d$ for $G$, with depth not larger than the depth of $\leq$, and with the additional property that for every $v\in V(G)$, if $C,C^{\prime}$ are distinct connected components of $G-V_{\preceq v}$ (or of $G$), then the vertices of $C$ and $C^{\prime}$ are incomparable with respect to $\preceq$.
Let $G$ be a graph. A graph $H$ with vertex set $\{v_{1},\ldots,v_{n}\}$ is a minor of $G$, written $H\preceq G$, if there are connected and pairwise vertex disjoint subgraphs $H_{1},\ldots,H_{n}\subseteq G$ such that if $\{v_{i},v_{j}\}\in E(H)$, then there are $w_{i}\in V(H_{i})$ and $w_{j}\in V(H_{j})$ such that $\{w_{i},w_{j}\}\in E(G)$. We call the subgraph $H_{i}$ the branch set of the vertex $v_{i}$ in $G$. If all $H_{i}$ have radius at most $r$, then we say that $H$ is a depth-$r$ minor of $G$, in symbols $H\preceq_{r}G$. If $G$ is a graph and =$\mathcal{H}=\{H_{1},\ldots,H_{n}\}$ is a set of pairwise vertex disjoint subgraphs of $G$, then the graph $H$ with vertex set $\{v_{1},\ldots,v_{n}\}$ and edges $\{v_{i},v_{j}\}\in E(H)$ if and only if there is an edge between a vertex of $H_{i}$ and a vertex of $H_{j}$ in $G$, the minor induced by $\mathcal{H}$. If $\bigcup_{1\leq i\leq n}V(H_{i})=V(G)$, we call $\mathcal{H}$ a minor model of $H$ that subsumes all vertices of $G$.
We denote by $K_{t}$ the complete graph on $t$ vertices. We denote by $G_{m\times n}$ the grid with $m$ rows and $n$ columns, that is, the graph with vertex set $\{v_{i,j}\mid 1\leq i\leq m,1\leq j\leq n\}$ and edges $\{v_{i,j},v_{i^{\prime},j^{\prime}}\}$ for $|i-i^{\prime}|+|j-j^{\prime}|=1$.
@@
Theorem 2.4 (Excluded Grid Theorem)
There exists a function $g$ such that for every integer $t\geq 1$, every graph of treewidth at least $g(t)$ contains the $t\times t$ grid as a minor. Furthermore, such a grid minor can be computed in polynomial time.
@@
Theorem 2.5 (Planar Excluded Grid Theorem, see Theorem 7.23 of [27])
@@
Theorem 2.6 (Courcelle’s Theorem [11])
@@ 3 The proof
For $q,k\in\mathbb{N}$, a graph $G$ is $(q,k)$-unbreakable if for every separation $(A,B)$ of $G$ of order at most $k$, we have $|A|\leq q$ or $|B|\leq q$.
Lemma 3.1
@@
Proof 3.2
Let $G=G_{m\times m}$ and let $A,B$ be a separation of order at most $k$ in $G$. We show that $|A|\leq k^{2}$ or $|B|\leq k^{2}$. The statement is trivially true if $m\leq k$ as then $|V(G)|\leq k^{2}$. Hence, assume that $m>k$, and let $\overline{z}\coloneqq z_{1},\ldots,z_{k}=A\cap B$.
We call a row or a column of $G$ untouched if it does not intersect $\overline{z}$ and touched if it does intersect $\overline{z}$. As there are at least $k+1$ rows (resp. columns) there is at least one untouched row (resp. column). An untouched row (resp. column) must be contained in one part of the separation. Assume w.l.o.g. that $A$ contains an untouched row. Then $A$ contains all untouched columns and, by symmetry, all untouched rows. Therefore every element in $B$ must be part of a touched row and a touched column, and, as there are at most $k$ touched rows resp. columns, there are at most $k^{2}$ such elements.
@@
Corollary 3.3 (of Lemma 3.1)
For all integers $k,m$ with $k<m$, for all connected graphs $G$ for which there exists a minor model $\mathcal{G}_{m,m}=\{H_{i,j}~{}|~{}1\leq i,j\leq m\}$ that subsumes all vertices of $G$, $k$ rounds of recursive elimination on $G$ can delete vertices in at most $k^{2}$ many different branch sets.
Definition 3.4
Let $G$ be a $K_{5}$-minor-free graph and let $k,d\in\mathbb{N}$. Assume there exists a minor model $\mathcal{G}_{m,m}=\{H_{i,j}\mid 1\leq i,j\leq m\}$ (for $m\geq 4k+7$) that subsumes all vertices of $G$ and that induces a supergraph of the grid $G_{m,m}$. We call the branch set $H_{i,j}$ $(k,d)$-safe if $32k+3\leq i,j\leq m-2k-$ 3 and if $H_{i^{\prime},j^{\prime}}$ contains no vertex of degree at least $d+1$ (in $G$) for $|i-i^{\prime}|,|j-j^{\prime}|\leq 2k+2$.
Lemma 3.5
Let $G$ be a $K_{5}$-minor-free graph and let $k,d\in\mathbb{N}$. Assume there exists a minor model $\mathcal{G}_{m,m}=\{H_{i,j}\mid 1\leq i,j\leq m\}$ (for $m\geq 4k+7$) that subsumes all vertices of $G$ and that induces a supergraph of the grid $G_{m,m}$. Assume $H_{i,j}$ is $(k,d)$-safe. Let $a\in V(H_{i,j})$, let $B\subseteq V(G)\setminus\{a\}$ with $|B|\leq k$ and let $x,y\in V(G)\setminus(B\cup\{a\})$ be of degree at least $d+1$. Then $x$ and $y$ are connected in $G-B$ if and only if $x$ and $y$ are connected in $G-B-a$.
Figure 1: Construction of a $K_{5}$ minor as soon as a branch set (here $H_{1}$) connected to a branch set (here $H_{2}$) that is at distance more than 2 in the grid [27, Figure 7.10]. The blue part marks the “outer part” of the border of thickness $2$, the “inner part” decomposes into a green, yellow and gray part.
Proof 3.6
@@
We define sets $X_{\ell}$ for $10\leq\ell\leq k+$ 1 of branch sets as follows. Let $X_{0}$ be the set consisting only of $H_{i,j}$ and for $\ell\geq 1$ let $X_{\ell}$ be the set of all $H_{i^{\prime},j^{\prime}}$ with $2\ell-1\leq|i-i^{\prime}|,|j-j^{\prime}|\leq 2\ell$ that do not already belong to $X_{\ell-1}$. The sets $X_{\ell}$ are the borders (of thickness $2$) of the $(4\ell+1)\times(4\ell+1)$-subgrid around $H_{i,j}$. Observe that the vertices $x$ and $y$ do not belong to any of the $X_{\ell}$, as $H_{i,j}$ is $(k,d)$-safe by assumption. For $10\leq\ell\leq k+$ 1 let $Y_{\ell}$ be the subgraph of $G$ induced by the vertices of $X_{\ell}$.
We claim that for $11\leq\ell\leq k+$ 1, the sets $Y_{\ell}$ are connected sets that separate $a$ from $x$ and analogously $a$ from $y$. Clearly, the $Y_{\ell}$ are connected. Now observe that there is no edge between a vertex of $\bigcup_{0\leq i\leq\ell-1}Y_{i}$ and a vertex of $G-\bigcup_{0\leq i\leq\ell}Y_{i}$. The existence of such a connection would create a $K_{5}$ minor, see [27, Figure 7.10] or Figure 1. Hence, any path between $a$ and $x$ (or $y$) must pass through $Y_{\ell}$.
As $|B|\leq k$, there is one $Y_{\ell}$ with $11\leq\ell\leq k+$ 1 that does not intersect $B$. Let $u$ be the first time that $P$ visits $Y_{\ell}$ on its way from $x$ to $a$ and let $v$ be the last time that $P$ visits $Y_{\ell}$ on its way from $a$ to $y$. As $Y_{\ell}$ is connected, we can reroute the subpath between $u$ and $v$ through $Y_{\ell}$ and thereby construct a path between $x$ and $y$ in $G-B-a$.
@@
Corollary 3.7
Let $G$ be a $K_{5}$-minor-free graph and let $k,d\in\mathbb{N}$. Assume there exists a minor model $\mathcal{G}_{m,m}=\{H_{i,j}\mid 1\leq i,j\leq m\}$ (for $m\geq 4k+7$) that subsumes all vertices of $G$ and that induces a supergraph of the grid $G_{m,m}$. Assume $H_{i,j}$ is $(k,d)$-safe. Then every vertex $a\in H_{i,j}$ is irrelevant, i.e., $G-a\in\mathscr{C}_{k,d}$ if and only if $G\in\mathscr{C}_{k,d}$.
Proof 3.8
Let $H\coloneqq G-a$. We have to prove that $H\in\mathscr{C}_{k,d}$ implies $G\in\mathscr{C}_{k,d}$. Hence, assume $H\in\mathscr{C}_{k,d}$. Let $\leq_{H}$ be an elimination order to degree $d$ of depth $k$ for $H$. We also assume that $\leq_{H}$ satisfies the property of Lemma 2.3, that is, for every $v\in V(G)$, if $C,C^{\prime}$ are distinct connected components of $G-V_{\leq_{H}v}$, then the vertices of $C$ and $C^{\prime}$ are incomparable with respect to $\leq_{H}$. Let $A_{0}$ be the connected component of $G$ containing $a$. Note that $A_{0}$ may break into multiple connected components in $H=G-a$. We show below that at most one of these connected components contains a vertex of degree excedeeing $d$.
For $1\leq i\leq k$, we define inductively:
  • $m_{i}$ as the unique $\leq_{H}$-minimal element of $A_{i-1}\setminus\{a\}$ (if it exists) such that there exists a vertex $v$ with $m_{i}\leq_{H}v$ and of degree at least $d+1$ in $H[A_{i-1}\setminus\{a\}]$. If there is no such element $m_{i}$, the process stops.
    Let us prove that there is at most one candidate for $m_{i}$. Assume that there are incomparable $m$ and $m^{\prime}$ satisfying these conditions. This means that there are vertices $v$ and $v^{\prime}$ of degree at least $d+1$ in $A_{i-1}\setminus\{a\}$ (hence of degree at least $d+1$ in $G$) with $m\leq_{H}v$ and $m^{\prime}\leq_{H}v^{\prime}$. Note that we have $m\not\leq_{H}v^{\prime}$ because $\leq_{H}$ is a tree order. By Lemma 2.3, we have that $v,v^{\prime}$ are both in $A_{i-1}\setminus\{a\}$, i.e. in the connected component of $a$ in $G-\{m_{1},\ldots,m_{i-1}\}$ (it follows also by induction that $\{m_{1},\ldots,m_{i-1}\}=V_{\leq_{H}m_{i-1}}$, hence we may apply the lemma). Hence, $v$ and $v^{\prime}$ are connected in $G-\{m_{1},\ldots,m_{i-1}\}$. With Lemma 3.5, we also have that $v$ and $v^{\prime}$ are connected in $G-\{a,m_{1},\ldots,m_{i-1}\}$.
    We take a witness path from $v$ to $v^{\prime}$. Since $m\leq_{H}v$ and $m\not\leq_{H}v^{\prime}$, this path must contain two adjacent vertices $w$, $w^{\prime}$ with $m\leq_{H}w$ and $m\not\leq_{H}w^{\prime}$. This contradicts the fact that $\leq_{H}$ is an elimination order satisfying the property of Lemma 2.3. Therefore, there is at most one possible such $m_{i}$.
  • $T_{i}:=\{v\in A_{i-1}~{}:~{}m_{i}\not\leq_{H}v\}$. These are the vertices that are disconnected from $A_{i}$ in $H$, but not in $G$. As explained above, they all have degree at most $d$.
@@
The processes stops after at most $k$ rounds. When the process stops, we have then defined $m_{i}$, $T_{i}$, and $A_{i}$ up to $i=\omega$, with $\omega\leq k$ and every element in $A_{\omega}$ has degree at most $d$ in $H[A_{\omega}]$.
We then define the new order $\leq_{G}$ as follows:
  • for all $x,y$ other than $a$ and that are not in any of the $T_{i}$ nor in $A_{\omega}$, we have $x\leq_{G}y$ if and only if $x\leq_{H}y$,
  • for all $x$ in $A_{\omega}\cup\bigcup\limits_{i\leq\omega}T_{i}$, we have $m_{i}\leq_{G}x$ for all $i\leq\omega$. Note that $a$ is in $A_{\omega}$.
Note that all the elements in $A_{\omega}$, and the $T_{i}$’s, together with $a$ are $\leq_{G}$-maximal.
We now prove that this new order is indeed an elimination order to degree $d$ of depth $k$ for $G$. We have that $\leq_{G}$ is a tree order and that it has depth at most $k$. Let us now take a vertex $b$ and study $S_{b}$. Recall the definition of $S_{b}$ from Definition 2.1. As we have two orders, we distinct $S^{G}_{b}$ from $S^{H}_{b}$.